OpenStudy (anonymous):

Garret is in the middle of a 20 minute bike ride. It takes 4 minutes for him to go a kilometer, so he'll end up going 5 kilometers. Suppose k is the number of kilometers Garret has already biked. Which expressions match below? Expressions: 5(4−k) 5−k 20−4k 4(5-k) 4k

OpenStudy (anonymous):

Is there anymore to the question?

OpenStudy (vijeya3):

@starwarsfan123 your question seems incomplete.What do you want to compute an expression for?

OpenStudy (javk):

if I were to ignore the first paragraph entirely then I would say that the expressions that match are `5-k` and `20-4k = 4(5-k)`. evidently because the latter is a multiple of the former. if however, you are trying to find an expression for how much longer he needs to travel then I would say that \[speed=\frac{ distance }{ time }\] speed is `1/4`, the distance covered up until this point is `5-k` and let the time remaining be called `t` in that case, just plug all the values in and you'll get \[\frac{ 1 }{ 4 } = \frac{ 5-k }{t }\] rearrang it and you'll get \[4(5-k)=t\]